# Asymptotic Normality of Subgraph Counts in Sparse Inhomogeneous Random Graphs

### v1

@@ 3. $(3)$
$Z(H,G_{n})=\frac{\Delta(H,G_{n})}{\sqrt{\mathrm{Var}[\Delta(H,G_{n})]}}\stackrel{{\scriptstyle D}}{{\rightarrow}}N(0,\,1),$ (2.11)
@@
  • $n\rho_{n}^{m_{1}(H)}\ll 1$: In this case, $\Delta_{2}(H,G_{n})$ has asymptotically negligible contribution to the total variance of $X(H,G_{n})$ and the asymptotic distribution of $X(H,G_{n})$ is determined entirely by the fluctuations of $\Delta_{1}(H,G_{n})$.
  • $n\rho_{n}^{m_{1}(H)}\gg 1$: In this case, $\Delta_{1}(H,G_{n})$ has asymptotically negligible contribution to the total variance of $X(H,G_{n})$ and the asymptotic distribution of $X(H,G_{n})$ is determined entirely by the fluctuations of $\Delta_{2}(H,G_{n})$.
  • $n\rho_{n}^{m_{1}(H)}\rightarrow c\in(0,\infty)$: In this case, both $\Delta_{1}(H,G_{n})$ and $\Delta_{2}(H,G_{n})$ contribute to the fluctuations of $X(H,G_{n})$. In particular, $(\Delta_{1}(H,G_{n}),\Delta_{2}(H,G_{n}))$ (after appropriate standardization) converges jointly independent normals with non-trivial variances.
@@
Recall that Theorem 2.11 is under the standing assumption that $n\rho_{n}^{m(H)}\ll 1$. By Proposition 2.9 $m(H)<m_{1}(H)$ for every subgraph $H$ with $|E(H)|\geq 1.$ Hence, all the above three regimes in Theorem 2.11 are relevant.
Another common consequence Theorem 2.11 in all the three cases, is that $Z(H,G_{n})$ is asymptotically normal. We know from Theorem 2.6 that the same is also true when $W$ is $H$-regular (recall (2.6)). Hence, we have the following corollary:
@@
$Z(H,G_{n})=\frac{\Delta(H,G_{n})}{\sqrt{\mathrm{Var}[\Delta(H,G_{n})]}}\stackrel{{\scriptstyle D}}{{\rightarrow}}N(0,\,1),$
where $Z(H,G_{n})$ is as defined in (2.2).
This shows that $Z(H,G_{n})$ is asymptotically normal (irrespective of whether or not $W$ is $H$-regular), for any $\rho_{n}\ll 1$ above the containment threshold. In previous works, this result has been established for either certain types of graph $H$ or in certain sub-regimes of sparsity above the containment threshold. Our result closes all the gaps in both these directions by showing that asymptotic normality of $Z(H,G_{n})$ holds for all fixed graph $H$ and for the full range of the sparsity parameter above the containment threshold. A detailed comparison with existing results is given in Section 2.4.2.
@@ 2.3. A General Conditional CLT and Proof Outline
Suppose Assumption 2.3 holds and $\frac{\mathrm{Var}[\Delta_{1}(H,\,G_{n})]}{\mathrm{Var}[\Delta(H,G_{n})]}\to\kappa$ for some constant $\kappa>0$. Then, for $Z\sim N(0,1)$,
@@
The proof of Proposition 2.15 is given in Section 4. The proof uses on Stein’s method based on dependency graphs, which bounds the distance to normality in the Wasserstein distance. Note that the Wasserstein distance in the LHS of (2.12) is itself random (since it is defined conditional on $\mathcal{F}(\bm{U}_{n})$), hence the convergence in (2.12) is shown to hold in probability. An important feature of result in Proposition 2.15 is that it holds regardless of whether or not the graphon $W$ is $H$-irregular, whenever $\Delta_{1}(H,G_{n})$ has a non-trivial contribution to the total variance. Hence, Proposition 2.15 can used to prove the assertions of Theorem 2.6 and Theorem 2.11 as follows:
  • $W$ is $H$-regular (Theorem 2.6): This is proved in Section 5. The first step is a variance computation to show that
@@
      Hence, only $\Delta_{2}(H,\,G_{n})$ contributes to the fluctuations of $\Delta(H,G_{n})$. In this case, $\Delta_{2}(H,\,G_{n})$ is $U$-statistic of order $|E(H)|$ which is not first-order degenerate $W$ is is not $H$-irregular (since $W$ is is not $H$-regular). Hence, invoking standard results about the asymptotic normality of $U$-statistics the result in (2.9) follows.
@@ 2.4.1. Erdős-Rényi Model
In the Erdős-Rényi model $G(n,\rho_{n})$ (which corresponds to $W\equiv 1$), it is well-known that $Z(H,G_{n})$ is asymptotically normal if and only if $n\rho_{n}^{m(H)}\gg 1$ and $n^{2}(1-\rho_{n})\gg 1$. This result was established independently by Ruciński [51] (using the method of moments), Barbour [3] (using Stein’s method), Nowicki [46] (using $U$-statistics), and Janson [33] (using martingales), among others (see [34, Chapter 6] for a comprehensive treatment). Note that in the sparse regime $\rho_{n}\ll 1$, the condition $n^{2}(1-\rho_{n})\gg 1$ is always satisfied, ensuring asymptotic normality for all subgraph counts $X(H,G_{n})$ whenever $n\rho_{n}^{m(H)}\gg 1$. Corollary 2.13 shows that the same result hold for the sparse $W$-random model as well.
@@ 2.4.2. Sparse Graphon Model
Central limit theorem for the subgraph counts in the sparse graphon model was first derived by Bickel et al. [8]. The sparsity conditions in their result depends on whether $H$ is acyclic (that is, $H$ does not contain a cycle) or whether it contains a cycle. Specifically, they showed the following results. Throughout, suppose $G_{n}\sim G(n,\rho_{n},W)$.
  • $H$ is acyclic: In this case, [8, Theorem 1] shows that $X(H,G_{n})$ (appropriately standardized) is asymptotically normal, whenever $n\rho_{n}\gg 1$. [^2] Technically, Bickel et al. [8] considers the number of induced copies, but their proof can be easily adapted for $X(H,G_{n})$. Note that for any acyclic graph $m(H)=1$. Hence, [8, Theorem 1] covers the full sparsity regime when $H$ is acyclic.
  • $H$ is acyclic: In this case, [8, Theorem 1] shows that $X(H,G_{n})$ is asymptotically normal, whenever $n\rho_{n}^{\frac{|V(H)|}{2}}\gg 1$. Recalling (2.1) note that
@@
    Hence, there is a gap between the sparsity regime covered by [8, Theorem 1] and the full sparsity regime ($n\rho_{n}^{m(H)}\gg 1$), when $H$ contains a cycle.
Later, Zhang and Xia [56] strengthened above the results by proving Edgeworth expansions and Berry-Essen bounds, under stronger sparsity conditions. Very recently, using the method of cumulants Liu and Privault [40] derived normal approximation bounds in the Kolmogorov distance for $Z(H,G_{n})$, when $H$ is strongly balanced. Notably, the rate of convergence in [40, Corollary 7.8] depends on whether $n\rho_{n}^{m_{1}(H)}\ll 1$ or $n\rho_{n}^{m_{1}(H)}\gg 1$, which, interestingly, coincides with the threshold identified in Theorem 2.11, where the dominant contribution to the total variance shifts from $\Delta_{1}(H,G_{n})$ and $\Delta_{2}(H,G_{n})$. The results in [40], however, do not consider the case where $H$ is not strongly balanced, or when $n\rho_{n}^{m_{1}(H)}\asymp 1$. In comparison, our results apply to all fixed subgraphs (not necessarily strongly balanced) and cover the entire sparsity regime, albeit without explicit rates of convergence.
@@ 2.4.3. Dense Graphon Model
Throughout this paper we have considered the sparse graphon model (where $\rho_{n}\ll 1$). In the dense regime (when $\rho_{n}=1$), the asymptotic normality of $X(H,G_{n})$ for all graphs $H$ was established by Féray et al. [26] (see also [2, 23, 35, 57] for alternate proofs, extensions, and related results). However, the limiting normal distribution of $X(H,G_{n})$ as derived in [26], can be degenerate depending on the structure of $W$. This degeneracy phenomenon was investigated by Hladkỳ et al. [28] when $H=K_{r}$ is the $r$-clique (the complete graph on $r$ vertices), for general graphs $H$ by Bhattacharya et al. [5], and for joint distributions by Chatterjee et al. [18]. Specifically, these results show that when $W$ is not $H$-regular (recall Definition 2.4), then $X(H,G_{n})$ is asymptotically Gaussian, with a normalization factor of $n^{|V(H)|-\frac{1}{2}}$. However, when $W$ is $H$-regular, then the normalization factor becomes $n^{|V(H)|-1}$ and the limiting distribution of $X(H,G_{n})$ has, in general, a Gaussian component and another independent (non-Gaussian) component which is a (possibly) infinite weighted sum of centered chi-squared random variables. This degeneracy phenomenon also appears in the subsequent work of Chatterjee and Huang [17] on the fluctuations of the largest eigenvalue. Recently, Huang et al. [32] established an invariance principle for $X(H,G_{n})$ that encompasses higher-order degeneracies.
@@
The proof of Proposition 2.2 is a standard application of the first and second moment methods. Specifically, it relies on the following lemma, which derives the asymptotic orders of the mean and variance of $X(H,G_{n})$.
@@
  • $H$ is balanced: In this case, $m(H)=\frac{|E(H)|}{|V(H)|}$ and $n\rho_{n}^{m(H)}\ll 1$ implies $n^{|V(H)|}\rho_{n}^{|E(H)|}\ll 1$, that is, $\mathbb{E}[X(H,G_{n})]\to 0$ (recall (3.1)). Therefore, by Markov’s inequality,
    $\mathbb{P}(X(H,G_{n})>0)$ $=\mathbb{P}(X(H,G_{n})\geq 1)\leq\mathbb{E}[X(H,G_{n})]\rightarrow 0.$
  • $H$ is unbalanced: Let $F\subseteq H$ be a subgraph of $H$ such that $m(H)=\frac{|E(F)|}{|V(F)|}$. Then $n\rho_{n}^{m(H)}\ll 1$ implies $n^{|V(F)|}\rho_{n}^{|E(F)|}\to 0$, that is, $\mathbb{E}[X(F,\,G_{n})]\to 0$. This implies,
    $\mathbb{P}(X(H,\,G_{n})>0)\leq\mathbb{P}(X(F,\,G_{n})>0)\leq\mathbb{E}[X(F,\,G_{n})]\to 0.$
@@
$\frac{\mathrm{Var}[X(H,\,G_{n})]}{(\mathbb{E}[X(H,\,G_{n})])^{2}}\lesssim_{H,W}\max_{F\subseteq H,\,|V(F)|\geq 1}\frac{1}{n^{|V(F)|}\rho_{n}^{|E(F)|}}.$ (3.3)
@@
$\frac{X(H,\,G_{n})}{\mathbb{E}[X(H,\,G_{n})]}\stackrel{{\scriptstyle P}}{{\rightarrow}}1\quad\implies\quad X(H,G_{n})\stackrel{{\scriptstyle P}}{{\rightarrow}}\infty,$
since $\mathbb{E}[X(H,\,G_{n})]\rightarrow\infty$, whenever $n\rho_{n}^{m(H)}\gg 1$. This completes the proof of Proposition 2.2. $\Box$
@@
The result in (3.1) is an immediate consequence of (1) $t(H,W)>0$ (recall Assumption 2.3).
We now proceed with the proof of (3.2). To this end, let $\mathcal{H}_{n}=\{H_{1},\,H_{2},\,\ldots,\,H_{M}\}$ be the collection of all copies of $H$ in the complete graph $K_{n}$. Here, $M=X(H,K_{n})$ the number of copies of $H$ in $K_{n}$. Now, define $I_{s}:=\bm{1}\{H_{s}\subseteq G_{n}\}$, for $1\leq s\leq M$. Then $\mathbb{E}[I_{s}]=\rho_{n}^{|E(H)|}t(H,\,W)$, for all $1\leq s\leq M.$ Furthermore,
$\mathrm{Var}[X(H,G_{n})]$
@@
$=\sum_{F\subseteq H:|V(F)|\geq 1}\sum_{\stackrel{{\scriptstyle H_{s},\,H_{t}\in\mathcal{H}_{n}}}{{H_{s}\cap H_{t}\simeq F}}}\mathrm{Cov}[I_{s},\,I_{t}]$
$=\sum_{F\subseteq H:|V(F)|\geq 1}\sum_{\stackrel{{\scriptstyle H_{s},\,H_{t}\in\mathcal{H}_{n}}}{{H_{s}\cap H_{t}\simeq F}}}[\rho_{n}^{2|E(H)|-|E(F)|}t(H_{s}\cup H_{t},\,W)-(\rho_{n}^{|E(H)|}t(H,\,W))^{2}].$ (3.4)
@@
Now, assume that $\rho_{n}\ll 1$. Fix a subgraph $F\subseteq H$ such that $|V(F)|\geq 1.$ By Lemma and fix copies $H_{s},\,H_{t}\in\mathcal{H}_{n}$ so that $H_{s}\cap H_{t}$ is isomorphic to $F.$ Then
$\mathrm{Cov}[I_{s},\,I_{t}]$ $\leq\mathbb{E}[I_{s}I_{t}]=\rho_{n}^{|E(H_{s}\cup H_{t})|}t(H_{s}\cup H_{t},\,W)$
[…]
Also, note that for a fixed subgraph $F\subseteq H,$ the number of ordered pairs $(H_{s},\,H_{t})$ with $H_{s}\cap H_{t}\simeq F$ is $O_{H}(n^{2|V(H)|-|V(F)|}).$ Hence, (3.4) can be expressed as
$\mathrm{Var}[X(H,G_{n})]$ $\lesssim_{H,W}\sum_{F\subseteq H:|V(F)|\geq 1}n^{2|V(H)|-|V(F)|}\rho_{n}^{2|E(H)|-|E(F)|}$
$\lesssim_{H,W}\max_{F\subseteq H:|V(F)|\geq 1}\frac{n^{2|V(H)|}\rho_{n}^{2|E(H)|}}{n^{|V(F)|}\rho_{n}^{|E(F)|}}.$ (3.6)
This proves the upper bound in (3.2)
To show the lower bound, we need the following lemma. The proof is given in Appendix A.2.

### v2

@@ 3. $(3)$
$Z_{n}(H,G_{n})=\frac{\Delta(H,G_{n})}{\sqrt{\mathrm{Var}[\Delta(H,G_{n})]}}\stackrel{{\scriptstyle D}}{{\rightarrow}}N(0,1),$ (2.11)
@@
  • $n\rho_{n}^{m_{1}(H)}\ll 1$: In this case, $\Delta_{2}(H,G_{n})$ has asymptotically negligible contribution to the total variance of $X_{n}(H,G_{n})$ and the asymptotic distribution of $X_{n}(H,G_{n})$ is determined entirely by the fluctuations of $\Delta_{1}(H,G_{n})$.
  • $n\rho_{n}^{m_{1}(H)}\gg 1$: In this case, $\Delta_{1}(H,G_{n})$ has asymptotically negligible contribution to the total variance of $X_{n}(H,G_{n})$ and the asymptotic distribution of $X_{n}(H,G_{n})$ is determined entirely by the fluctuations of $\Delta_{2}(H,G_{n})$.
  • $n\rho_{n}^{m_{1}(H)}\rightarrow c\in(0,\infty)$: In this case, both $\Delta_{1}(H,G_{n})$ and $\Delta_{2}(H,G_{n})$ contribute to the fluctuations of $X_{n}(H,G_{n})$. In particular, $(\Delta_{1}(H,G_{n}),\Delta_{2}(H,G_{n}))$ (after appropriate standardization) converges jointly to independent normals with non-trivial variances.
@@
Recall that Theorem 2.11 is under the standing assumption that $n\rho_{n}^{m(H)}\ll 1$. By Proposition, we know that 2.9 $m(H)<m_{1}(H)$, for every subgraph $H$ with $|E(H)|\geq 1.$ Hence, all the above three regimes in Theorem 2.11 are relevant.
Another consequence Theorem 2.11 is that $Z_{n}(H,G_{n})$ is asymptotically normal in all the three cases. We know from Theorem 2.6 that the same is also true when $W$ is $H$-regular (recall (2.6)). Hence, we have the following corollary:
@@
$Z_{n}(H,G_{n})=\frac{\Delta(H,G_{n})}{\sqrt{\mathrm{Var}[\Delta(H,G_{n})]}}\stackrel{{\scriptstyle D}}{{\rightarrow}}N(0,1),$
where $Z_{n}(H,G_{n})$ is as defined in (2.2).
This shows that $Z_{n}(H,G_{n})$ is asymptotically normal (irrespective of whether or not $W$ is $H$-regular), for any $\rho_{n}\ll 1$ above the containment threshold. In previous works, this result has been established for either certain types of graph $H$ or in certain sub-regimes of sparsity above the containment threshold. Our result closes all the gaps in both these directions by showing that asymptotic normality of $Z_{n}(H,G_{n})$ holds for all fixed graph $H$ and for the full range of the sparsity parameter above the containment threshold. A detailed comparison with existing results is given in Section 2.4.2.
@@ 2.3. A General Conditional CLT and Proof Outline
Suppose Assumption 2.3 holds. Then, for $Z\sim N(0,1)$,
@@
The proof of Proposition 2.15 is given in Section 4. The proof uses on Stein’s method based on dependency graphs, which bounds the distance to normality in the Wasserstein distance. Note that the Wasserstein distance in the LHS of (2.12) is itself random (since it is defined conditional on $\mathcal{F}(\bm{U}_{n})$), hence the convergence in (2.12) is shown to hold in probability. An important feature of result in Proposition 2.15 is that it holds regardless of whether or not the graphon $W$ is $H$-regular. Hence, Proposition 2.15 can be used to prove the assertions of Theorem 2.6 and Theorem 2.11 as follows:
  • $W$ is $H$-regular (Theorem 2.6): This is proved in Section 5. The first step is a variance computation showing that
@@
      Hence, only $\Delta_{2}(H,G_{n})$ contributes to the fluctuations of $\Delta(H,G_{n})$. In this case, $\Delta_{2}(H,G_{n})$ is a non-degenerate $U$-statistic of order $|E(H)|$ (since $W$ is not $H$-regular). Hence, invoking standard results about the asymptotic normality of $U$-statistics the result in (2.9) follows.
@@ 2.4.1. Erdős-Rényi Model
In the Erdős-Rényi model $G(n,\rho_{n})$ (which corresponds to $W\equiv 1$), it is well-known that $Z_{n}(H,G_{n})$ is asymptotically normal if and only if $n\rho_{n}^{m(H)}\gg 1$ and $n^{2}(1-\rho_{n})\gg 1$. This result was established independently by Ruciński (1988) (using the method of moments), Barbour et al. (1989) (using Stein’s method), Nowicki (1989) (using $U$-statistics), and Janson (1990) (using martingales), among others (see Janson et al. (2011, Chapter 6) for a comprehensive discussion). Note that in the sparse regime $\rho_{n}\ll 1$, the condition $n^{2}(1-\rho_{n})\gg 1$ is always satisfied, ensuring asymptotic normality for all subgraph counts $X_{n}(H,G_{n})$ whenever $n\rho_{n}^{m(H)}\gg 1$. Corollary 2.13 shows that the same result hold for the sparse $W$-random model as well.
@@ 2.4.2. Sparse Graphon Model
Central limit theorem for subgraph counts in the sparse graphon model was first derived by Bickel et al. (2011). The sparsity conditions in their result depends on whether $H$ is acyclic (that is, $H$ does not contain a cycle) or whether it contains a cycle. Specifically, they showed the following results. Throughout, suppose $G_{n}\sim G(n,\rho_{n},W)$.
  • $H$ is acyclic: In this case, Bickel et al. (2011, Theorem 1) shows that $X_{n}(H,G_{n})$ (appropriately standardized) is asymptotically normal, whenever $n\rho_{n}\gg 1$. [^2] Technically, Bickel et al. (2011) considers the number of induced copies, but their proof can be easily adapted for $X_{n}(H,G_{n})$. Note that for any acyclic graph $m(H)<1$. Hence, there is a gap between the sparsity regime covered in Bickel et al. (2011, Theorem 1) and the full sparsity regime ($n\rho_{n}^{m(H)}\gg 1$), when $H$ is acyclic.
  • $H$ is not acyclic: In this case, Bickel et al. (2011, Theorem 1) shows that $X_{n}(H,G_{n})$ is asymptotically normal, whenever $n\rho_{n}^{\frac{|V(H)|}{2}}\gg 1$. Recalling (2.1) note that
@@
    Therefore, there is again a gap between the sparsity regime in Bickel et al. (2011, Theorem 1) and the full sparsity regime ($n\rho_{n}^{m(H)}\gg 1$), when $H$ contains a cycle.
Later, Zhang and Xia (2022) strengthened above the results by proving Edgeworth expansions and Berry-Essen bounds, under stronger sparsity conditions. Very recently, using the method of cumulants Liu and Privault (2026) derived normal approximation bounds in the Kolmogorov distance for $Z_{n}(H,G_{n})$, when $H$ is strongly balanced. Notably, the rate of convergence in Liu and Privault (2026, Corollary 7.8) depends on whether $n\rho_{n}^{m_{1}(H)}\ll 1$ or $n\rho_{n}^{m_{1}(H)}\gg 1$, which, interestingly, coincides with the threshold identified in Theorem 2.11, where the dominant contribution to the total variance shifts from $\Delta_{1}(H,G_{n})$ and $\Delta_{2}(H,G_{n})$. The results in Liu and Privault (2026), however, do not consider the case where $H$ is not strongly balanced, or when $n\rho_{n}^{m_{1}(H)}\asymp 1$. In comparison, our results apply to all fixed subgraphs (not necessarily strongly balanced) and cover the entire sparsity regime, albeit without explicit rates of convergence.
@@ 2.4.3. Dense Graphon Model
Throughout this paper we have considered the sparse graphon model (where $\rho_{n}\ll 1$). In the dense regime (when $\rho_{n}\asymp 1$), the asymptotic normality of $X_{n}(H,G_{n})$ for all graphs $H$ was established by Féray et al. (2020) (see also Austern and Orbanz (2022), Delmas et al. (2021), Kaur and Röllin (2021), Zhang (2022) for alternate proofs, extensions, and related results). However, the limiting normal distribution of $X_{n}(H,G_{n})$ as derived in Féray et al. (2020), can be degenerate depending on the structure of $W$. This degeneracy phenomenon was investigated by Hladkỳ et al. (2021) when $H=K_{r}$ is the $r$-clique (the complete graph on $r$ vertices), for general graphs $H$ by Bhattacharya et al. (2023), and for joint distributions by Chatterjee et al. (2024). Specifically, these results show that when $W$ is not $H$-regular (recall Definition 2.4), then $X_{n}(H,G_{n})$ is asymptotically Gaussian, with a normalization factor of $n^{|V(H)|-\frac{1}{2}}$. However, when $W$ is $H$-regular, then the normalization factor becomes $n^{|V(H)|-1}$ and the limiting distribution of $X_{n}(H,G_{n})$ has, in general, a Gaussian component and another independent (non-Gaussian) component which is a (possibly) infinite weighted sum of centered chi-squared random variables. This degeneracy phenomenon also appears in the subsequent work of Chatterjee and Huang (2024) on the fluctuations of the largest eigenvalue. Recently, Huang et al. (2024) established an invariance principle for $X_{n}(H,G_{n})$ that encompasses higher-order degeneracies.
@@
The proof of Proposition 2.2 is a standard application of the first and second moment methods. Specifically, it relies on the following lemma, which derives the asymptotic orders of the mean and variance of $X_{n}(H,G_{n})$.
@@
  • $H$ is balanced: In this case, $m(H)=\frac{|E(H)|}{|V(H)|}$ and $n\rho_{n}^{m(H)}\ll 1$ implies $n^{|V(H)|}\rho_{n}^{|E(H)|}\ll 1$, that is, $\mathbb{E}[X_{n}(H,G_{n})]\to 0$ (recall (3.1)). Therefore, by Markov’s inequality,
    $\mathbb{P}(X_{n}(H,G_{n})>0)$ $=\mathbb{P}(X_{n}(H,G_{n})\geq 1)\leq\mathbb{E}[X_{n}(H,G_{n})]\rightarrow 0.$
  • $H$ is unbalanced: Let $F\subseteq H$ be a subgraph of $H$ such that $m(H)=\frac{|E(F)|}{|V(F)|}$. Then $n\rho_{n}^{m(H)}\ll 1$ implies $n^{|V(F)|}\rho_{n}^{|E(F)|}\to 0$, that is, $\mathbb{E}[X_{n}(F,G_{n})]\to 0$. This implies,
    $\mathbb{P}(X_{n}(H,G_{n})>0)\leq\mathbb{P}(X_{n}(F,G_{n})>0)\leq\mathbb{E}[X_{n}(F,G_{n})]\to 0.$
@@
$\frac{\mathrm{Var}[X_{n}(H,G_{n})]}{(\mathbb{E}[X_{n}(H,G_{n})])^{2}}\lesssim_{H,W}\max_{F\subseteq H,\,|V(F)|\geq 1}\frac{1}{n^{|V(F)|}\rho_{n}^{|E(F)|}}.$ (3.5)
@@
$\frac{X_{n}(H,G_{n})}{\mathbb{E}[X_{n}(H,G_{n})]}\stackrel{{\scriptstyle P}}{{\rightarrow}}1\quad\implies\quad X_{n}(H,G_{n})\stackrel{{\scriptstyle P}}{{\rightarrow}}\infty,$
since $\mathbb{E}[X_{n}(H,G_{n})]\rightarrow\infty$, whenever $n\rho_{n}^{m(H)}\gg 1$. This completes the proof of Proposition 2.2. $\Box$
@@
The result in (3.1) is an immediate consequence of (1) and the condition $t(H,W)>0$ from Assumption 2.3.
We now proceed with the proof of (3.4). To this end, let $\mathcal{H}_{n}=\{H_{1},H_{2},\ldots,H_{M}\}$ be the collection of all copies of $H$ in the complete graph $K_{n}$. Here, $M=X_{n}(H,K_{n})$ is the number of copies of $H$ in $K_{n}$. Now, define $I_{s}:=\bm{1}\{H_{s}\subseteq G_{n}\}$, for $1\leq s\leq M$. Then $\mathbb{E}[I_{s}]=\rho_{n}^{|E(H)|}t(H,W)$, for all $1\leq s\leq M.$ Furthermore,
$\mathrm{Var}[X_{n}(H,G_{n})]$
@@
$=\sum_{F\subseteq H:|V(F)|\geq 1}\sum_{\begin{subarray}{c}H_{s},H_{t}\in\mathcal{H}_{n}\\ H_{s}\cap H_{t}\simeq F\end{subarray}}\mathrm{Cov}[I_{s},I_{t}]$
$=\sum_{F\subseteq H:|V(F)|\geq 1}\sum_{\begin{subarray}{c}H_{s},H_{t}\in\mathcal{H}_{n}\\ H_{s}\cap H_{t}\simeq F\end{subarray}}[\rho_{n}^{2|E(H)|-|E(F)|}t(H_{s}\cup H_{t},W)-(\rho_{n}^{|E(H)|}t(H,W))^{2}],$ (3.6)
@@
To show the upper bound and lower bound statements in (3.4), we need the following lemma, which essentially states that for each subgraph $F\subseteq H,$ one can find a suitable number of copies $H_{s},H_{t}$ of $H$ such that $H_{s}\cap H_{t}\simeq F$ and $H_{s}\cup H_{t}$ has positive homomorphism density $t(H_{s}\cup H_{t},W)>0$ with respect to $W.$ The lemma is proved in Appendix A.2.